# latexFont shape

### v1

@@ 4.2 Overall Architecture
Reference Branch. For textual expressions, the common representations of their pointing objects are obtained by directly sending them to the text encoder of CLIP [12]. The channels of the resulting representations are transformed to $c_{d}$ to achieve a better trade-off between efficiency and performance, and we denote them as $F^{cate}\in\mathbb{R}^{c_{d}}$.
@@
The pipeline of extracting common representation from referring images is cascaded by a foreground prediction network based on the encoder-decoder structure and a masked average pooling (MAP) function. In particular, the former is used to generate the visual features and foreground predictions of the referring images, and the latter is employed to filter out irrelevant information in the referring images. By default, the ICON [56] model with ResNet-50 [83] backbone is selected as our foreground prediction network.
@@
Given $K$ referring images with spatial size $H\times W$ in a certain category, the visual features with spatial size $\frac{H}{32}\times\frac{W}{32}$ and foreground maps with spatial size $H\times W$ are first extracted from the encoder and decoder of the foreground prediction network. And we denote them as $\{F^{ref}_{k}\}_{k=1}^{K}$ and $\{M^{ref}_{k}\}_{k=1}^{K}$, respectively.
Then, $\{F^{ref}_{k}\}_{k=1}^{K}$ and $\{M^{ref}_{k}\}_{k=1}^{K}$ are fed into the MAP function to calculate representations associated with the foreground objects. This calculation is defined as follows:
@@
$E^{obj}_{k}=\mathcal{F}\mathrm{{}_{conv1\times 1}}\bigg{(}\frac{\sum_{2d}(\mathcal{F}\mathrm{{}_{down}}(M^{ref}_{k})\otimes F^{ref}_{k})}{\sum_{2d}(F^{ref}_{k})}\bigg{)},$ (1)
@@
where $\otimes$ represents the element-wise multiplication operation, ()$\mathcal{F}\mathrm{{}_{down}}(\cdot)$ is a bilinear downsampling operation for shape matching, $\sum_{2d}(\cdot)$ accumulates the feature values along the spatial dimension, and ()$\mathcal{F}\mathrm{{}_{conv1\times 1}}(\cdot)$ is a $1\times 1$ convolution that transforms the channel of its input to $c_{d}$.
Finally, the common representations of this category, i.e., $E^{cate}$, is attained by averaging these object representations.
@@
Segmentation Branch. Our segmentation branch is built upon an encoder-decoder structure as well, which has been widely used in COD research. Note that as the goal of this paper is to provide a new scheme for directionally segmenting the camouflaged objects, we do not put too much effort on the architectural design. In fact, we find that under the proposed scheme, even a simple segmentation network can still perform well. Thus, ResNet-50 [83] is adopted as the encoder, and the features from its last three layers are selected as visual representations, following previous works [1, 3]. And the decoder is a convolutional head consisting of two convolutional layers for the identification of camouflaged objects. We also propose two new modules, namely referring mask generation (RMG) and referring feature enrichment (RFE), which are added between the encoder and the decoder, to take advantage of the common representations from the reference branch for the explicit segmentation of camouflaged targets.
@@
Given an image with spatial size $H\times W$, we first extract multi-scale features from the last three stages of the encoder. And the channels of these features are converted to $c_{d}$ using $1\times 1$ convolutions. We denote these features as $\{F^{seg}_{j}\}_{j=2}^{4}$, where $F^{seg}_{j}\in\mathbb{R}^{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\times(c_{d}+8)}$.
Then, the resulting features are fed into the RMG module together with the common representations from the reference branch to generate fusion features and a referring mask, which are denoted as $F^{f}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times c_{d}}$ and $\mathcal{H}^{m}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 1}$. This process can be defined as:
$F^{f},\mathcal{H}^{m}=\mathrm{RMG}(\{F^{seg}_{j}\}_{j=2}^{4},F^{cate}).$ (2)
@@
Next, the fusion features are enriched with the guidance of the prior mask to highlight the camouflaged targets in the RFE module, which can be defined as:
$E^{enr}=\mathrm{RFE}(F^{f},\mathcal{H}^{m}).$ (3)
Finally, the resulting features are fed into the decoder to generate the segmentation map $M^{seg}\in\mathbb{R}^{H\times W\times 1}$.
@@ 4.3 Referring Mask Generation
In order to identify the camouflaged objects accurately from their highly similar surroundings according to the image guidance, a pixel-level comparison between the common representations and the visual features should be performed. However, the appearance of a camouflaged object may be significantly different from that of the referring objects even if they belong to the same category. In addition, the common representations and the visual features come from different information sources. Such a large information difference may interfere with the comparison process, making the localization process of the camouflaged targets difficult.
Figure 7: Structural details of the proposed Referring Mask Generation (RMG) module. Note that ‘C’ is the concatenation operation, ‘AT’ is the affine transformation, ‘CLSTM’ is the convolutional LSTM, and ’Dconv’ is the dynamic convolution.
Motivated by the recent works on multimodal fusion [84, 70], we solve this issue by performing a dual-source information fusion (DSF) between the common representations and the visual features, as illustrated in Fig. 7. To promote the interaction of these two kinds of information, the spatial position information is first injected into the visual features. Specifically, each position of the visual features is concatenated with an 8d embedding vector akin to the implementation of  [85]. We denote the visual features equipped with spatial information as $\{x^{seg}_{j}\}_{j=2}^{4}$, where $x^{seg}_{j}\in\mathbb{R}^{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\times(c_{d}+8)}$. Then, we apply an affine transformation on the visual features under the guidance of the common representations. In particular, two linear layers are utilized to map the common representations to two coefficient vectors, which are applied on the visual features followed by convolution and ReLU operations. The obtained results are denoted as $\{y_{j}\}_{j=2}^{4}$, where $y_{j}\in\mathbb{R}^{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\times c_{d}}$. And this process is defined as follows:
@@
$\mathcal{\gamma}_{j}=\mathcal{F}_{mlp1}(F^{cate}),\mathcal{\beta}_{j}=\mathcal{F}_{mlp2}(F^{cate}),$ (5)
@@
Figure 8: Structural details of the proposed Referring Feature Enrichment (RFE) module. Note that ‘R’ is the resize operation, ‘$1\times 1$’ is a convolutional layer with kernel size 1, and ‘Conv Block’ is composed of two convolutional layers with kernel size 3.
Furthermore, multi-scale fusions are performed on the dual-source fusion features from the three levels via convolutional LSTM [86] to enhance the robustness of our method to targets with different scales. This process can be formulated as follows:
@@
where ()$\mathcal{F}\mathrm{{}_{up}}(\cdot)$ is a bilinear upsampling operation for shape matching and $\mathcal{F}_{clstm}$ denotes a cell of convolutional LSTM. Note that $h_{4}=c_{4}=y_{4}$, and the last hidden state, namely $h_{2}\in\mathbb{R}^{\frac{H}{8}\times\frac{H}{8}\times c_{d}}$, is used as the fused visual feature $F^{f}$.
TABLE I: Comparison of the COD models with their Ref-COD counterparts. ‘R-50’: ResNet-50 [83], ‘E-B4’: EfficientNet-B4 [87], ‘R2-50’: Res2Net-50 [88], ‘$\rm R^{3}$-50’: Triple ResNet-50 [2]. ‘Ours’: R2CNet. ‘-RefS’: the model with image references composed of salient objects. ‘-RefT’: the model with text references. ‘Attribute’: the attribute of each network, ‘Single-obj’: the scene of a single camouflaged object, ‘Multi-obj’: the scene of multiple camouflaged objects, ‘Overall’: all scenes containing camouflaged objects. $\uparrow$’: the higher the better, $\downarrow$’: the lower the better.
@@
DGNet-RefS E-B4 20.10 7.24 95.06 0.821 0.891 0.696 0.032 0.827 0.890 0.703 0.031 0.748 0.879 0.607 0.045
DGNet-RefT E-B4 20.04 7.24 94.74 0.824 0.891 0.701 0.032 0.830 0.892 0.709 0.031 0.745 0.873 0.596 0.046
@@
Figure 9: PR and $F_{\beta}$ curves of the COD methods and their Ref-COD variants. The standard COD methods are shown via blue solid lines. The Ref-COD methods with image references and text references are shown via blue solid lines and black dashed lines, respectively.
Finally, the common representations are compared with the representations of each position of the fusion features to produce a referring mask. Inspired by [89], this target indexing process is realized through a dynamic convolution. In particular, the common representations and the fusion features are utilized as a reference-guided dynamic kernel and the input, respectively.
@@ 4.4 Referring Feature Enrichment
To be specific, the prior mask and the fusion features are first resized to the shapes of the aforementioned three features, i.e., $\{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\}_{j=2}^{4}$, respectively. Then, the resized masks and features at the same scale are fused by concatenation, and the output features of different scales, termed as $\{F^{scale}_{j}\}_{j=2}^{4}$, where $F^{scale}_{j}\in\mathbb{R}^{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\times c_{d}}$, are also concatenated together to enrich the capability of identifying camouflaged objects. We denote the enriched feature as $F^{enr}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times c_{d}}$. As mentioned in PFENet [69], tiny objects may become unclear in the down-sampled feature maps. Thus, we build a similar path that crosses scales from the finer feature to the coarse one to achieve the interaction between them. In addition, we also apply supervision on $\{F^{scale}_{j}\}_{j=2}^{4}$ to make the yielded feature $E^{enr}$ more robust, and the corresponding maps is denoted as $\{M^{scale}_{j}\}_{j=2}^{4}$, where $M^{scale}_{j}\in\mathbb{R}^{H\times W\times 1}$.
@@ 5.1 Experiment Setup
Training/Testing Protocols.
Considering that the goal of Ref-COD is to generate a binary foreground map closing to the annotation, the structure loss [90] consisting of BCE loss and IoU loss, which has been widely used in many binary segmentation tasks, can be adopted as the optimization objectives of the Ref-COD models. Specifically, this function is formulated as:
@@ 5.2 Quantitative Evaluation
Comparison with Baselines. To verify the effectiveness of our R2CNet, we first compare it with its baseline variant, which is a generic COD model based on the proposed encoder-decoder architecture. In this baseline model, the last three FPN features from its encoder are fused via the up-bottom path, and the fused feature is fed to its decoder to directly segment camouflaged objects without any reference. As shown in Tab. I, our R2CNet surpasses the baseline model by a large margin across all metrics in the overall scene (i.e., test split) of R2C7K. To evaluate the performance of the model more comprehensively, we further divide the overall scene into two groups, namely the scene of single camouflaged objects and the scene of multiple camouflaged objects. It can be observed that the R2CNet outperforms its COD counterpart in all settings, which demonstrates that the referring images contribute to the identification of the camouflaged objects from the confusing scenes.

### v3

@@ 4.2 Overall Architecture
Reference Branch. The pipeline of extracting common representation from referring images is cascaded by a SOD network based on the encoder-decoder structure and a masked average pooling (MAP) function. In particular, the former is used to generate the visual features and foreground predictions of the referring images, and the latter one is employed to filter out irrelevant information in the referring images. By default, the pre-trained ICON [56] model with ResNet-50 [84] backbone is selected as our SOD network.
@@
Given $K$ referring images with salient target objects and spatial size $H\times W$, the visual features with spatial size $\frac{H}{32}\times\frac{W}{32}$ and foreground maps with spatial size $H\times W$ are first extracted from the encoder and decoder of the SOD network. We denote them as $\{F^{ref}_{k}\}_{k=1}^{K}$ and $\{M^{ref}_{k}\}_{k=1}^{K}$, respectively. Then, $\{F^{ref}_{k}\}_{k=1}^{K}$ and $\{M^{ref}_{k}\}_{k=1}^{K}$ are fed into the MAP function to calculate representations associated with the foreground objects, denoted as $F^{obj}_{k}\in\mathbb{R}^{c_{d}\times 1\times 1}$. This calculation can be formulated as:
@@
$F^{obj}_{k}=\mathcal{F}\mathrm{{}_{conv1\times 1}}\bigg{(}\frac{\sum_{2d}(\mathcal{F}\mathrm{{}_{down}}(M^{ref}_{k})\otimes F^{ref}_{k})}{\sum_{2d}(F^{ref}_{k})}\bigg{)},$ (1)
@@
where $\otimes$ is the element-wise multiplication operation, ()$\mathcal{F}\mathrm{{}_{down}}(\cdot)$ is a bilinear downsampling operation for shape matching, $\sum_{2d}(\cdot)$ accumulates the feature values along the spatial dimension, and ()$\mathcal{F}\mathrm{{}_{conv1\times 1}}(\cdot)$ is a $1\times 1$ convolution that transforms the channel of its input to $c_{d}$ to achieve a better trade-off between efficiency and performance. Finally, we abstract the common representations of target objects in the embedding space, denoted as $E\in\mathbb{R}^{c_{d}\times 1\times 1}$, by averaging these object representations.
@@
Segmentation Branch. Our segmentation branch is built upon an encoder-decoder structure as well, which has been widely used in COD research. Note that as the goal of this paper is to provide a new scheme for directionally segmenting the camouflaged objects, we do not put too much effort on the architectural design. In fact, we find that under the proposed scheme, even a simple segmentation network can still perform well. Thus, ResNet-50 [84] is adopted as the encoder, and the features from its last three layers are selected as visual representations, following previous works [1, 3]. The decoder is a convolutional head consisting of two convolutional layers for the identification of camouflaged objects. We also propose two new modules, namely referring mask generation (RMG) and referring feature enrichment (RFE), which are added between the encoder and the decoder, to take advantage of the common representations from the reference branch for the explicit segmentation of camouflaged targets.
@@
Given an image with camouflaged objects and spatial size $H\times W$, we first extract multi-scale features from the last three stages from the encoder. The channels of these features are converted to $c_{d}$ using $1\times 1$ convolutions. We denote these features as $\{F^{seg}_{j}\}_{j=2}^{4}$, where $F^{seg}_{j}\in\mathbb{R}^{c_{d}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$. Then, the resulting features are fed into the RMG module together with the common representations from the reference branch to generate fusion features and a referring mask, which are denoted as $F^{f}\in\mathbb{R}^{c_{d}\times\frac{H}{8}\times\frac{W}{8}}$ and $\mathcal{H}^{m}\in\mathbb{R}^{1\times\frac{H}{8}\times\frac{W}{8}}$. This process can be defined as:
$F^{f},\mathcal{H}^{m}=\mathrm{RMG}(\{F^{seg}_{j}\}_{j=2}^{4},E).$ (2)
@@
Next, the fusion features are enriched at multiple scales with the guidance of the prior mask to highlight the camouflaged targets in the RFE module, which is defined as:
$F^{enr}=\mathrm{RFE}(F^{f},\mathcal{H}^{m}).$ (3)
Finally, the resulting features are fed into the decoder to generate the segmentation map $M^{seg}\in\mathbb{R}^{1\times H\times W}$.
@@ 4.3 Referring Mask Generation
In order to identify the camouflaged objects accurately from their highly similar surroundings according to the image guidance, a pixel-level comparison between the common representations of target objects and the visual features should be performed. However, the appearance of a camouflaged object may be significantly different from that of the referring objects even if they belong to the same category. In addition, the common representations and the visual features come from different information sources. Such a large information difference may interfere with the comparison process, making the localization process of the camouflaged targets difficult.
Figure 7: Structural details of the proposed Referring Mask Generation (RMG) module. Note that ‘C’ is the concatenation operation, ‘AT’ is the affine transformation, ‘CLSTM’ is the convolutional long short-term memory, and ’Dconv’ is the dynamic convolution.
Motivated by recent works on multimodal fusion [85, 72], we solve this issue by performing a dual-source information fusion (DSF) between the common representations and the visual features, as illustrated in Fig. 7. To promote the interaction of these two kinds of information, the spatial position information is first injected into the visual features. Specifically, each position of the visual features is concatenated with an 8d embedding vector akin to the implementation of  [86]. We denote the visual features equipped with spatial information as $\{x^{seg}_{j}\}_{j=2}^{4}$, where $x^{seg}_{j}\in\mathbb{R}^{(c_{d}+8)\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$. Then, we apply an affine transformation on the visual features under the guidance of the common representations. In particular, two linear layers are utilized to map the common representations to two coefficient vectors, which are applied on the visual features followed by convolution and ReLU operations. The obtained results are denoted as $\{y_{j}\}_{j=2}^{4}$, where $y_{j}\in\mathbb{R}^{c_{d}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$. This process is defined as follows:
@@
$\mathcal{\gamma}_{j}=\mathcal{F}_{mlp1}(E),\mathcal{\beta}_{j}=\mathcal{F}_{mlp2}(E),$ (5)
@@
Figure 8: Structural details of our Referring Feature Enrichment (RFE) module. Note that ‘R’ is the resize operation, ‘$1\times 1$’ is a convolutional layer with kernel size 1, and ‘Conv Block’ is composed of two convolutional layers with kernel size 3.
Furthermore, multi-scale fusions (MSF) are performed on the dual-source fusion features along the up-bottom path via convolutional LSTM [87] to enhance the robustness of our method to targets with different scales. This process can be formulated as follows:
@@
where ()$\mathcal{F}\mathrm{{}_{up}}(\cdot)$ is a bilinear upsampling operation for shape matching and $\mathcal{F}_{clstm}$ denotes a cell of convolutional LSTM. Note that the initial state $h_{4}=c_{4}=y_{4}$, and the last hidden state, i.e., $h_{2}\in\mathbb{R}^{c_{d}\times\frac{H}{8}\times\frac{H}{8}}$, is used as the fused feature $F^{f}$.
TABLE II: Comparisons of popular COD models with their Ref-COD counterparts. All models are evaluated on a NVIDIA RTX 3090 GPU. ‘R-50’: ResNet-50 [84], ‘E-B4’: EfficientNet-B4 [88], ‘R2-50’: Res2Net-50 [89], ‘$\rm R^{3}$-50’: Triple ResNet-50 [2], ‘Swin-S’: SwinTransformer-S [90], ‘-Ref’: models with image references composed of salient objects, ‘Attribute’: attributes of each network, ‘Single-obj’: scenes of a single camouflaged object, ‘Multi-obj’: scenes of multiple camouflaged objects, ‘Overall’: all scenes containing camouflaged objects, $\uparrow$’: the higher the better, $\downarrow$’: the lower the better.
@@
DGNet-Ref E-B4 20.10 7.24 95.06 0.821 0.891 0.696 0.032 0.827 0.890 0.703 0.031 0.748 0.879 0.607 0.045
$\rm VSCode_{2024}$ [91] Swin-S 74.72 59.81 76.81 0.819 0.879 0.702 0.033 0.825 0.880 0.706 0.032 0.750 0.868 0.651 0.043
VSCode-Ref Swin-S 76.63 64.28 65.26 0.832 0.891 0.714 0.030 0.838 0.892 0.718 0.029 0.766 0.880 0.662 0.041
$\rm ZoomNext_{2024}$ [92] $\rm R^{3}$-50 28.46 185.79 66.29 0.838 0.897 0.742 0.032 0.843 0.898 0.750 0.031 0.777 0.880 0.655 0.040
ZoomNext-Ref $\rm R^{3}$-50 30.32 197.43 52.81 0.850 0.909 0.755 0.027 0.859 0.910 0.762 0.026 0.788 0.892 0.675 0.037
@@
Figure 9: PR and $F_{\beta}$ curves of the COD methods and their Ref-COD variants. The results of standard COD methods and their Ref-COD counterparts are illustrated via blue dashed lines and red solid lines respectively.
Finally, the common representations are compared with the representations of each position of the fusion features to produce a referring mask, namely $\mathcal{H}^{m}$. Inspired by [93, 94], this target-matching (TM) process is realized through a dynamic convolution operation. In particular, the common representations and the fusion features are utilized as reference-guided dynamic kernel and input, respectively.
@@ 4.4 Referring Feature Enrichment
To be specific, the prior mask and the fusion features are first resized to the shapes of the aforementioned three features, i.e., $\{\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}\}_{j=2}^{4}$, respectively. Then, the resized masks and features at the same scale are fused by concatenation, and the output features of different scales, termed as $\{F^{scale}_{j}\}_{j=2}^{4}$, where $F^{scale}_{j}\in\mathbb{R}^{c_{d}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$, are also concatenated together to enrich the capability of identifying camouflaged objects. We denote the enriched feature as $F^{enr}\in\mathbb{R}^{c_{d}\times\frac{H}{8}\times\frac{W}{8}}$. As mentioned in PFENet [70], tiny objects may become unclear in the down-sampled feature maps. Thus, we build a similar path that crosses scales from the finer feature to the coarse one to achieve the interaction between them. In addition, we also apply supervision on $\{F^{scale}_{j}\}_{j=2}^{4}$ to make the yielded feature $F^{enr}$ more robust, and the corresponding maps is denoted as $\{M^{scale}_{j}\}_{j=2}^{4}$, where $M^{scale}_{j}\in\mathbb{R}^{1\times H\times W}$.
@@ 5 Experiments
In this section, we first introduce the experiment settings of this paper, i.e., training & testing protocols, hyperparameter details, and evaluation metrics. Then, we conduct quantitative comparisons between the Ref-COD methods and the corresponding COD methods. Particularly, we compare the proposed R2CNet with the baseline model on different camouflaged scenes. we also apply the design of Ref-COD on 7 recent state-of-the-art  COD methods to verify its generality. Next, we report ablative study results to investigate the effectiveness of each component and our design choice. Finally, we carry out visualization comparisons to gain more intuitive insight into our Ref-COD.
@@ 5.1 Experiment Setup
Training & Testing Protocols. Considering that the goal of Ref-COD is to generate a binary foreground map closing to the annotation, the structure loss [95] consisting of BCE loss and IoU loss, which has been widely used in many binary segmentation tasks, can be adopted as the optimization objectives of the Ref-COD models. Specifically, this function can be formulated as:
@@
Metrics. Following the standard protocol in COD, we adopt four common metrics for evaluation, including mean absolute error (M) [14], structure-measure (Sm[15], adaptive E-measure ($\alpha$E) [16] and weighted F-measure ($w$F) [17]. To be specific, M measures the absolute difference between the predicted mask and ground-truth (GT); Sm measures the region-aware and object-aware structural similarity between the predicted map and GT; $\alpha$E measures the element-wise and image-level similarity. $w$F is an exhaustive measure of both recall and precision between the prediction and GT. In addition, we draw the precision-recall (PR) curves and $F_{\beta}$-threshold ($F_{\beta}$) curves for further comparison.
TABLE III: Application of the paradigm of our Ref-COD to recent published salient object detection models.
@@ 5.2 Quantitative Evaluation
Comparison with Baselines. To verify the effectiveness of our R2CNet, we first compare it with its baseline variant, which is a standard COD model based on encoder-decoder architecture. In this baseline model, the last three features from its encoder are fused along the up-bottom path, following the fashion in FPN, and the fused feature is fed to its decoder to directly segment camouflaged objects without any reference. As shown in Tab. II, our R2CNet surpasses the baseline model by a large margin across all metrics in the overall scene (i.e., test split) of R2C7K. To evaluate the performance of the model more comprehensively, we further divide the overall scene into two groups, namely the scene of single camouflaged objects and the scene of multiple camouflaged objects. It can be observed that our R2CNet still outperforms its COD baseline in all settings, which demonstrates that the referring images contribute to the identification of the camouflaged objects from the confusing scenes.
Application to Existing COD Methods. To verify the generality of our Ref-COD design, we also apply this idea to existing COD methods. These methods are chosen according to three criteria: a) recently published, b) representative, and c) with open source code. In particular, the methods to participate in the adaptation here include: PFNet [31], PreyNet [24], SINetV2 [1], BSANet [37], BGNet [4], ZoomNet [2], DGNet [3], VSCode [91], and ZoomNext [92]. As shown from the third to the last row of Tab. II and Fig. 9, the Ref-COD variants of these models achieve better performance than their original versions in all metrics. It should be noted that our lightweight baseline model even achieves a performance comparable to recent COD methods after incorporating reference images. This indicates that our method does not rely on powerful segmentation networks. The introduction of the reference branch can help significantly improve the performance of a COD model even though it is not strong, reflecting the effectiveness of our Ref-COD idea. Moreover, the relative order of metrics values for Ref-COD methods is almost consistent with that of their COD counterparts, demonstrating that the distinct advantages of these COD algorithms can still be preserved when applying our design to them.
Application to Existing SOD Methods. To further verify the generality of our Ref-COD design on Salient Object Detection (SOD) task, we also conduct experiments on the CoSOD3K [99] dataset to apply the design of our Ref-COD on the latest SOD models, i.e., VSCode [91] and VST++ [100]. As shown in Tab. III, the Ref-SOD variants achieve better performance across all metrics compared to the original SOD methods.
5.3 Ablative Studies
Number of referring images. We ablate the number of referring images, i.e. $K$, for its impact on the Ref-COD. Considering that there are 5 referring images in each category of the test split of the R2C7K dataset, the value of $K$ changes from 0 to 5. When ‘$50<K<$ 5’, the performance of our R2CNet is calculated by averaging three evaluations, in which this model is guided by $K$ referring images randomly sampled from a certain category of the test split. The experimental results in Fig. 10 show that the performance of our R2CNet improves with the increase of the number of referring images. We believe this is because the acquired common information is less affected by sample differences during this process, and thus a more accurate location of camouflaged targets is achieved.